# ROCKET-2: Steering Visuomotor Policy via Cross-View Goal Alignment

### v1

@@ 4 Experiments
We aim to address the following questions: (1) How does ROCKET-2 perform in terms of both accuracy and efficiency during inference? (2) Can ROCKET-2 follow the intention of a human from a cross-camera view? (3) Under what circumstances does ROCKET-2 fail to work? (4) How important are landmarks in cross-view goal alignment? (5) Can ROCKET-2 interpret goal views from cross-episode scenarios? (6) Which modules contribute effectively to training ROCKET-2?
@@ 4.1 Experimental Setup
Implementation Details
@@
We present the model architecture, hyperparameters, and optimizer configurations of ROCKET-2 in Table 1. During training, each trajectory is divided into segments of length 128 to reduce memory requirements. We initialize the view backbone that is used to encode RGB images with DINO weights and freeze it for training efficiency. During inference, ROCKET-2 can access up to 128 key-value attention caches of past observations. Most training parameters follow those from prior works such as Cai et al. (2024b); Baker et al. (2022).
Figure 3: The Evaluation Metric is Spatial-Sensitive. $\checkmark$ and $\times$ indicate the correct and incorrect objects for interaction, respectively. None of the task configurations were seen during the training.
@@
Figure 4: Performance-Efficiency Comparison on the Minecraft Interaction Benchmark. The x-axis represents inference speed (FPS), and the y-axis shows the interaction success rate. Numbers in parentheses indicate the Molmo invocation interval, where larger values mean higher FPS. “+ track” denotes real-time SAM-2 segmentation between Molmo calls, increasing inference time (applicable only to ROCKET-1). In most cases, ROCKET-2 achieves 3×3$\times$3 × to 6×6$\times$6 × faster while matching or surpassing ROCKET-1’s peak performance.
Environment and Benchmark
We use the original Minecraft 1.16.5 (Guss et al., 2019; Lin et al., 2023; Cai et al., 2024a) as our testing environment, which accepts mouse and keyboard inputs and outputs a $640\times 360$ RGB image at each step. Following Cai et al. (2024b), we employ the Minecraft Interaction Benchmark to evaluate the agent’s interaction capabilities. This benchmark includes six categories and a total of 12 tasks, covering all basic Minecraft interaction types: Hunt, Mine, Interact, Navigate, Tool, and Place. As this benchmark emphasizes object interaction and spatial localization, its evaluation criteria are more stringent than those in Lifshitz et al. (2023) and Cai et al. (2023). We show some examples in Figure 3. In the “hunt the sheep in the right fence” task, success requires the agent to kill the sheep within the right fence, while killing it in the left fence results in failure. Similarly, in the “place the oak door on the diamond block” task, success is only achieved if the oak door is adjacent to the diamond block on at least one side.
Baselines
We compare our ROCKET-2 with the following instruction-following baselines: (1) STEVE-1 (Lifshitz et al., 2023): An instruction-following agent fine-tuned from VPT (Baker et al., 2022), capable of solving various short-horizon tasks. We use the text-conditioned version of STEVE-1 for comparison. (2) GROOT-1 (Cai et al., 2023): A reference-video-conditioned policy designed for open-ended tasks, trained on 2,000 hours of long-form videos using latent variable models. (3) ROCKET-1 (Cai et al., 2024b): A segmentation-conditioned policy capable of mastering 12 interaction tasks. While it achieves a high interaction success rate, its reliance on SAM-2’s real-time tracking during inference creates an efficiency bottleneck.
4.2 Performance-Efficiency Analysis
@@
Table 2: ROCKET-Series Inference Pipeline Details. Molmo can pinpoint the target object based on the task prompt. SAM uses the point to generate object mask $m_{t}$ w.r.t. $o_{t}$ and supports real-time object tracking.
@@
Figure 5: Case Study of Human-Agent Interaction. We demonstrate how a human interacts with ROCKET-2, leveraging its spatial reasoning abilities. (Top Row) The human specifies a hay bale (Refer to caption) that is not visible to ROCKET-2. By exploring the area around the visible landmark (house), ROCKET-2 successfully locates the goal. (Bottom Row) The human specifies a target tree in the presence of a tree distractor. ROCKET-2 accurately identifies the correct tree by reasoning about spatial relationships and landmarks. The agent’s trajectories are visualized in bird’s-eye view maps.
We demonstrate that cross-view goal specification (implemented as ROCKET-2) significantly improves inference speed while maintaining high interaction success. Following Cai et al. (2024b), we evaluate STEVE-1, GROOT-1, and variants of ROCKET-1 and our ROCKET-2 on the Minecraft Interaction Benchmark. Due to the limited spatial reasoning ability, STEVE-1 and GROOT-1 achieve success rates below $20\%$ on most tasks. We illustrate the inference pipeline of ROCKET-series agents in Table 2. Automated evaluation of ROCKETs relies on Molmo (Deitke et al., 2024) and SAM (Ravi et al., 2024) to generate a segmentation mask for the target object in the given views. In general, ROCKET-1 requires object masks for all agent observations, whereas ROCKET-2 only needs one or a few object masks. We observe that ROCKET-1 attains over $80\%$ success with high-frequency Molmo point corrections (every 3 frames) but suffers from slow inference. Lowering Molmo’s frequency greatly degrades ROCKET-1’s performance, with tasks like “collecting lava” failing entirely at 30-frame intervals. While one can enable SAM’s tracking mode to provide dense goal signals, it remains computationally expensive. In contrast, our ROCKET-2 decouples the goal specification from the agent view, it does not need frequent segmentation mask modification and can autonomously track the target object. It achieves comparable or superior performance to ROCKET-1 with a 3×3$\times$3 × to 6×6$\times$6 × inference speedup.
4.3 Intuitive Human-Agent Interaction
In Figure 5, we present two case studies illustrating ROCKET-2 interprets human intent under the cross-view goal specification interface. The first case (top row) involves a task requiring the agent to approach a hay bale ([Uncaptioned image]) located behind a house ([Uncaptioned image]). From the human view, both the house and the hay bale are visible, whereas ROCKET-2 initially observes only the house. A key challenge arises from the differing camera views: the human and ROCKET-2 perceive the scene from opposite sides of the house. To analyze the agent’s behavior, we visualize both its camera views and its trajectories on a bird’s-eye map. We observe that ROCKET-2 effectively infers the hay bale’s potential location and successfully navigates toward it. This is reflected in the increasing target visibility score and the movement of the predicted point. Interestingly, the bird’s-eye view reveals that ROCKET-2 approaches the target from both sides of the house, demonstrating diversity in route selection. The second case (bottom row) showcases ROCKET-2 ’s ability to distinguish between a distractor and the human-specified goal object, despite their visual similarity. This finding highlights that ROCKET-2 ’s spatial reasoning extends beyond object appearance and incorporates scene alignment for goal inference.
Figure 6: Visualization Analysis of Cross-View Alignment. The vision patches (identified by white grid) represent a chosen background landmark in the agent’s current view (instead of the goal object). We generated an attention map with the spatial fusion transformer using these patches as queries and the goal view patches as keys and values. We found that ROCKET-2 perfectly aligned with the selected landmarks across views.
4.4 Analyzing Failure Cases
We analyze failure cases in ROCKET-2 ’s task execution and identify three main issues: (1) Prediction Drift: When pursuing distant targets for extended periods, the predicted point gradually shifts away from the object. Since distant targets rely on temporal consistency from memory for recognition, but the model was only trained with memory lengths up to 128, it faces long-sequence generalization challenges during inference. (2) Distance Perception Error: When the goal and agent camera views differ significantly, the agent sometimes stops one step before reaching the target, leading to interaction failure. We observe that updating the goal view to the agent’s current observation alleviates this issue, likely due to the greater cross-view discrepancy encountered during inference compared to training. (3) Action Jitter: When inferring the original version of ROCKET-2, we observe significant action jitter, this could cause failures in precise interactions such as placing blocks. We find that incorporating previous actions during training and inference greatly improves action smoothness.
4.5 Landmarks Attention Visualization
Specifically, we prepare a current view observation and a third view with goal segmentation. Before being fed into the spatial transformer, both views are encoded into $14\times 14=196$ tokens: $\{\hat{o}_{t}^{i}\}_{i=1}^{196}$ and $\{h_{g}^{i}\}_{i=1}^{196}$ (notations are consistent with Sec. 3). We inspect the softmax-normalized attention map of the first self-attention layer in the spatial transformer, denoted as $\{a_{i,j}\}_{i,j=1}^{392}$, where $a_{i,197:392}$ represents the attention map generated by using patch $i$ from the current view as the query and all patches from the third view as keys and values. This map is overlaid on the third view (goal view) to reflect its responsiveness to patch $i$ in the current view. Since landmarks may span multiple patches, we aggregate the response maps of different patches to form the final attention map $\{m_{i}\}_{i=1}^{196}$:
$m_{i}=\frac{1}{|L|}\sum_{x\in L}a_{x,i+196},$ (6)
where $L$ denotes the set of patches in the current view representing a specific landmark. Notably, the selected landmarks do not overlap with the goal segmentation. As shown in Figure 6, we present four sets of data covering villages, plains, deserts, and forest terrains. In the first plot, the white grid indicates the selected landmark patches, while the third plot shows the third view response to the chosen landmarks. Our findings reveal that ROCKET-2 effectively matches cross-view consistency even under significant geometric deformations and distance variations. Surprisingly, in data point (4), even subtle forest depressions at a considerable distance are accurately matched.
Figure 7: Cross-Episode Generalization. The goal view does not exist within the agent’s world but originates from a different episode. We observe that the agent attempts to infer the semantic information underlying the goal specification.
4.6 Cross-Episode Goal Alignment
4.7 Ablation Studies on Auxiliary Objectives
Table 3: Ablation Study on Auxiliary Objectives. The final loss function for each row is the cumulative sum of all loss functions from the preceding rows.
Training Variants [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Avg. behavior cloning $0.52$ $0.78$ $0.65$ $0.65$ +target visibility $0.63$ $0.83$ $0.68$ $0.71$ +cross-view consistency $\mathbf{0.85}$ $\mathbf{0.97}$ $\mathbf{1.00}$ $\mathbf{0.94}$
To evaluate the impact of auxiliary losses on model performance, we define three variants: (1) only behavior cloning loss, (2) + target visibility loss, and (3) the full version with + cross-view consistency loss. We conduct experiments on three tasks: Navigate to House in a Village([Uncaptioned image]), Mine Emerald([Uncaptioned image]), and Interact with the Left Chest([Uncaptioned image]). We find that the BC-only variant achieves an average success rate of only $65\%$, demonstrating that the action signal is insufficient for learning spatial alignment. Adding target visibility loss improves performance by $6\%$, while further incorporating cross-view consistency loss boosts the success rate to $94\%$. This proves that leveraging temporal consistency and introducing vision-based auxiliary losses can greatly enhance cross-view goal alignment and inference-time decision-making capabilities.
5 Conclusions and Limitations
To improve human-agent interaction in embodied worlds, we propose a cross-view goal specification approach. Since behavior cloning alone fails to align the agent with human views, we introduce cross-view consistency and target visibility losses to enhance alignment. ROCKET-2 achieves state-of-the-art performance on the Minecraft Interaction Benchmark with a 3×3$\times$3 × to 6×6$\times$6 × efficiency boost. Visualizations and case studies validate our method. We also observe that ROCKET-2 struggles with visual reasoning when the discrepancy between the agent’s and human’s views is large. As memory length increases, the predicted points exhibit noticeable drift. We attribute this to the relabeled dataset, whose memory window is short and view variation is limited. Enhancing data quality could help address this issue.
6 Acknowledgements
This work is funded by the National Science and Technology Major Project #2022ZD0114902.

### v2

@@ 4 Experiments
We aim to address the following questions: (1) How does ROCKET-2 perform in terms of both accuracy and efficiency during inference? (2) Can ROCKET-2 follow the intent of a human from a cross-camera view in Minecraft? (3) Can ROCKET-2 adapt and generalize to other 3D environments? (4) How important are landmarks in cross-view goal alignment for ROCKET-2? (5) Can ROCKET-2 interpret goal views from cross-episode scenarios? (6) Which modules contribute effectively to training ROCKET-2?
@@ 4.1 Experimental Setup
We include all the details, such as hyperparameters, benchmarks, case studies and extensive experiments, in the Appendix.
Figure 5: Inference Pipelines on Minecraft Interaction Benchmark. GROOT-1 requires retreving a short video as its condition. ROCKET-1’s segmentation is coupled with the agent’s real-time observations, whereas ROCKET-2’s segmentation is solely tied to the third-person view.
Environment and Benchmark
We use Minecraft v1.16.5 (Guss et al., 2019; Cai et al., 2024) as the testing environment, which accepts mouse and keyboard inputs and returns a $640\times 360$ RGB image per step. Following Cai et al. (2025a), we employ the Minecraft Interaction Benchmark to evaluate the agent’s interaction capabilities. This benchmark includes six categories and a total of 12 tasks, covering all basic Minecraft interaction types: Hunt, Mine, Interact, Navigate, Tool, and Place. As this benchmark emphasizes object interaction and spatial localization, its evaluation criteria are more stringent than those in Lifshitz et al. (2023) and Cai et al. (2023). We present three examples in Figure 3, more can be found in appendix. In the “hunt the sheep in the right fence” task, success requires the agent to kill the sheep within the right fence, while killing sheep in the left fence results in failure.
Baselines
We compare our ROCKET-2 with the following instruction-following baselines: (1) STEVE-1 (Lifshitz et al., 2023): A text-conditioned agent fine-tuned from VPT (Baker et al., 2022), capable of solving various short-horizon tasks. (2) GROOT-1 (Cai et al., 2023): A video-conditioned policy designed for open-ended tasks, implemented with a VAE network. (3) ROCKET-1 (Cai et al., 2025a): A mask-conditioned policy capable of mastering 12 interaction tasks. Taking the hunt right sheep task as an example, we present the inference pipelines in Figure 5, where Molmo (Deitke et al., 2024) and SAM (Ravi et al., 2024) are combined to perform text-based instance segmentation. Given that ROCKET-2 relies on a third-person viewpoint for goal specification, in the Minecraft Interaction Benchmark, we set its third-person view to the observation acquired by the agent at the initial step. This view is then largely maintained, undergoing resets only at fixed periods (e.g., every 90 steps). This setup ensures that ROCKET-2 requires no additional privileged information.
Figure 6: Case Study of Human-Agent Interaction. We show how a human interacts with ROCKET-2, leveraging its spatial reasoning abilities. (Top Row) The human specifies a hay bale (Refer to caption) that is not visible to ROCKET-2. By exploring the area around the visible landmark (house), ROCKET-2 successfully locates the goal. (Bottom Row) Human specifies a target tree in the presence of a tree distractor. ROCKET-2 accurately identifies the correct tree by reasoning about spatial relationships. The trajectories are visualized in BEV maps.
4.2 Performance-Efficiency Analysis
Figure 4 presents our Minecraft Interaction Benchmark results, showcasing various agent configurations and their performance. Specifically, ROCKET-1 (10) performs object segmentation with Molmo + SAM every 10 steps, providing no masks for the intermediate 9 steps; its "+ track" variant leverages SAM’s more computationally efficient object tracking to causally generate masks for these frames. In contrast, ROCKET-2 utilizes cross-view segmentation, only requiring a single third-person view segmentation, with ROCKET-2 (90) resetting this view every 90 steps. Our initial observations highlight that STEVE-1 and GROOT-1 exhibit success rates below $20\%$ across most tasks, primarily due to their instructions’ limited spatial sensitivity. While ROCKET-1 achieves over $80\%$ success with high-frequency Molmo (every 3 frames), it suffers from slow inference; lowering Molmo’s frequency or enabling SAM tracking (though still expensive) significantly degrades its performance. Conversely, our ROCKET-2 agent, by decoupling goal specification from the agent’s current view, autonomously tracks targets without frequent mask modifications, achieving comparable or superior performance to ROCKET-1 with a remarkable 3×3$\times$3 × to 6×6$\times$6 × inference speedup.
4.3 Intuitive Human-Agent Interaction
In Figure 6, we present two case studies illustrating ROCKET-2 interprets human intent under the cross-view goal specification interface. The first case (top row) involves a task requiring the agent to approach a hay bale ([Uncaptioned image]) located behind a house ([Uncaptioned image]). From the human view, both the house and the hay bale are visible, whereas ROCKET-2 initially observes only the house. A key challenge arises from the differing camera views: the human and ROCKET-2 perceive the scene from opposite sides of the house. To analyze the agent’s behavior, we visualize both its camera views and its trajectories on a bird’s-eye map. We observe that ROCKET-2 effectively infers the hay bale’s potential location and successfully navigates toward it. This is reflected in the increasing target visibility score and the movement of the predicted point. Interestingly, the bird’s-eye view reveals that ROCKET-2 approaches the target from both sides of the house, demonstrating diversity in route selection. The second case (bottom row) showcases ROCKET-2 ’s ability to distinguish between a distractor and the human-specified goal object, despite their visual similarity. It highlights that agent’s spatial reasoning extends beyond object appearance.
Figure 7: Zero-Shot 3D Environment Generalization. ROCKET-2 can find and transport the victim to the stretcher in the unseen Unreal Engine. We map the left click/right click actions of Minecraft to the carry up/put down actions of Unreal to interact with objects, respectively.
4.4 Zero-Shot 3D Worlds Generalization
We demonstrate that by mapping the action space of Minecraft to unseen 3D games (refer to Table 1), ROCKET-2 achieves zero-shot generalization, despite being trained solely on the Minecraft dataset. We attribute this to two key factors: (1) the DINO-pretrained ViT backbone is frozen during training, preserving its general 3D view perception capability; and (2) the learned cross-view goal alignment skill is domain-agnostic. We evaluated ROCKET-2 in the ATEC 2025 AI and Robotics Challenge  [^1] https://github.com/atecup/atec2025_software_algorithm, which is built in the Unreal Engine and assigns agents to locate injured people and transport them to stretchers. The benchmark provides both textual (a long text description) and visual cues (a third-person view) about the locations of the victims. Note that, ROCKET-2 takes in solely visual cues (Figure 7). As shown in Table 2, ROCKET-2 outperforms the strong baseline provided by the organizers, a Gemma-3 and YOLO-based two-tiered decision system using both textual and visual cues, by $12\%$, despite not being fine-tuned on the Unreal. Furthermore, the Unreal environment provides observations at a $640\times 480$ resolution, a notable deviation from the $640\times 360$ resolution of the training dataset. This robustly substantiates that cross-view goal alignment enables the acquisition of transferable 3D decision representations within one environment and their successful migration to another. Such capabilities underscore the significant feasibility of developing general-purpose multi-task agents for diverse 3D environments.
Table 1: Bridging the Minecraft Action Space and Other 3D Games. “/” denotes the masked action.
Minecraft DeepMind Lab Doom Unreal $\text{forward}=1$ $a[3]=1$ $\text{discrete}[2]=1$ $\text{velocity}=+100$ $\text{back}=1$ $a[3]=-1$ $\text{discrete}[3]=1$ $\text{velocity}=-100$ $\text{attack}=1$ $a[4]=1$ $\text{discrete}[1]=1$ $\text{pick}=1$cdots $\cdots$ $\cdots$ $\cdots$ $\text{yaw}=x$ $a[0]=4.75x$ $\text{continuous}=x$ $\text{angular}=x$ $\text{pitch}=x$ $a[1]=2.78x$ / $\text{viewport}=x$
Table 2: Results on ATEC 2025 Robotics Challenge.
Agent ROCKET-2 (Zero-Shot) YOLO-Gemma 3 Rank $\mathbf{1/57}$ $20/57$ Success Rate $\bm{38}\%$ $26\%$
Figure 8: Visualization Analysis of Landmarks’ Alignment. The vision patches (identified by white grid) represent a chosen background landmark in the agent’s current view (instead of the goal object). We generate an attention map with the spatial fusion transformer using these patches as queries and the human view patches as keys and values. We find that ROCKET-2 well aligns selected landmarks across views.
4.5 Ablation Studies on Auxiliary Objectives
To evaluate the impact of auxiliary losses on model performance, we define three variants: (1) only behavior cloning loss, (2) + target visibility loss, and (3) the full version with + cross-view consistency loss. We conduct experiments on three tasks: Navigate to House in a Village([Uncaptioned image]), Mine Emerald([Uncaptioned image]), and Interact with the Left Chest([Uncaptioned image]). We find that the BC-only variant achieves an average success rate of only $65\%$, demonstrating that the action signal is insufficient for learning spatial alignment. Adding target visibility loss improves performance by $6\%$, while further incorporating cross-view consistency loss boosts the success rate to $94\%$. This proves that leveraging temporal consistency and introducing vision-based auxiliary losses can greatly enhance cross-view goal alignment and inference-time decision-making.
Table 3: Ablation Study on Auxiliary Objectives. The final loss function for each row is the cumulative sum of all loss functions from the preceding ones.
Model Variants [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Avg. behavior cloning $0.52$ $0.78$ $0.65$ $0.65$ + target visibility $0.63$ $0.83$ $0.68$ $0.71$ + cross-view consistency $\mathbf{0.85}$ $\mathbf{0.97}$ $\mathbf{1.00}$ $\mathbf{0.94}$
4.6 Landmarks Attention Visualization
Specifically, we prepare a current view observation and a third view with goal segmentation. Before being fed into the spatial transformer, both views are encoded into $14\times 14=196$ tokens: $\{\hat{o}_{t}^{i}\}_{i=1}^{196}$ and $\{h_{g}^{i}\}_{i=1}^{196}$ (notations are consistent with Sec. 3). We inspect the softmax-normalized attention map of the first self-attention layer in the spatial transformer, denoted as $\{a_{i,j}\}_{i,j=1}^{392}$, where $a_{i,197:392}$ represents the attention map generated by using patch $i$ from the current view as the query and all patches from the third view as keys and values. This map is overlaid on the third view (goal view) to reflect its responsiveness to patch $i$ in the current view. Since landmarks may span multiple patches, we aggregate the response maps of different patches to form the final attention map $\{m_{i}\}_{i=1}^{196}$: $m_{i}=\frac{1}{|L|}\sum_{x\in L}a_{x,i+196}$, where $L$ denotes the set of patches in the current view representing a specific landmark. Notably, the selected landmarks do not overlap with the goal segmentation. As shown in Figure 8, we present three sets of data covering villages, plains, and forest terrains. In the left plot, the white grid indicates the selected landmark patches, while the right plot shows the human-view attention response to the chosen landmarks. Our findings reveal that ROCKET-2 effectively matches cross-view consistency even under significant geometric deformations and distance variations. Surprisingly, in the last data point, even subtle forest depressions are accurately matched.
Figure 9: Cross-Episode Generalization. The goal view does not exist within the agent’s world but originates from a different episode. ROCKET-2 attempts to infer the semantic intent underlying the goal specification (building a bridge and approaching a house).
4.7 Cross-Episode Cross-View Goal Alignment
5 Conclusion
[…]
Acknoledgements
This work was supported by the National Science and Technology Major Project #2022ZD0114902. We sincerely appreciate their generous support, which enabled us to conduct this research.
Appendix
Implementation Details and Extended Results
Figure 10: Minecraft Demonstrations. ROCKET-2 showcases a diverse range of skills, including combat with monsters, locating the Ender portal, bridge construction, ore mining, and crafting. Cross-view perspectives with segmented and circled masks are displayed in the top right window. Predicted bounding boxes, visibility, and actions are overlaid on the images, highlighting the agent’s robust 3D visual goal understanding and alignment with human-specified targets.
Appendix A Environments
ROCKET-2 is trained in Minecraft (Cai et al., 2024) and evaluated not only within Minecraft but also across a diverse set of out-of-domain environments, including Unreal Engine (Zhong et al., 2024), Doom (Kempka et al., 2016), DeepMind Lab (Beattie et al., 2016), and several commercial Steam games. Appendix A provides an overview of their observation and action spaces. To enable cross-environment testing, we construct a rule-based mapping from Minecraft’s action space to those of the target environments (table 1), without relying on environment-specific fine-tuning or carefully crafted adaptations.
We present qualitative demonstrations in both Minecraft and unseen environments in Figure 10 and Figure 12, where cross-view images with human-provided masks serve as goal specifications. In Minecraft, ROCKET-2 showcases advanced interactive abilities, including defeating the Ender Dragon and constructing long-span bridges—tasks previously unattainable by existing agents. Remarkably, ROCKET-2 also exhibits strong generalization to unfamiliar environments: it performs visually-guided exploration in Unreal Engine, effectively transfers combat behaviors to Doom, and remains robust despite drastic differences in observation styles across games such as Portal and Dark Souls. These results highlight the model’s ability to leverage visual cues and make meaningful decisions in a zero-shot setting, underscoring the versatility and generality of cross-view goal alignment.
Table 4: Overview of Environments. The blue line separates the training environment from the zero-shot evaluation environments. Minecraft is our primary training environment, while others serve for zero-shot generalization assessment.
{tblr}
colspec = Q[m, l, wd=3.2cm, font=] X[l, m] X[l, m] , row1 = bg=tableheadergray, fg=black, font=, row2 = bg=trainingenvcolor, roweven[4-Z] = bg=tableheadergray, cell2-Z1 = cmd=, measure = vbox, Environment Observation Space Details Action Space Details
Minecraft (Training)
(Cai et al., 2024)
  • Visual: First-person RGB images ($640\times 360$).
  • Numerical: Comprehensive player information (e.g., health, food levels); not utilized during training and inference.
  • Discrete actions: attack, back, forward, hotbar, inventory, jump, left, right, sneak, sprint and use.
  • Continuous actions: adjusting pitch and yaw in $[-180,180]$.
  • For the output of agent, we take the hierarchical action space (Baker et al. (2022)) of 8,641 discrete button combinations and 121 camera movement bins.
[tablelineblue] Unreal Engine
(Zhong et al., 2024)
  • Visual: First-person RGB images as well as goal images ($640\times 480$). The observation resolution notably differs from that of Minecraft.
  • Goal: The agent must rescue objects and transport them to designated stretchers, requiring exploration of complex 3D environments guided by visual goal cues.
  • Discrete actions: carry, drop, jump, crouch, and open the door.
  • Continuous actions: adjusting angular velocity (turning), linear velocity (moving forward/backward), and viewpoint (masked).
ViZDoom
(Kempka et al., 2016)
  • Visual: First-person RGB view (e.g., $320\times 240$).
  • We use the basic config for demonstration.
  • 3 primary actions are enabled in the basic setting: move left/right, shoot (attack).
DeepMind Lab
(Beattie et al., 2016)
  • Visual: First-person RGB images ($640\times 480$).
  • We choose map seekavoid_arena_01, a fruit-gathering environment with navigation scenarios.
  • Discrete actions: move back/forward, fire, jump, crouch, and strafe left/right.
  • Continuous actions: adjusting angular velocity (turning), and viewpoint (masked) in angular per second of game time.
Steam Games
(Portal/Dark Souls)
  • Visual: First-person RGB view, resized to $640\times 360$.
  • Same as Minecraft, where output actions are mapped to human keyboard and mouse movements.
Appendix B Implementation Details
@@
We present the model architecture, hyperparameters, and optimizer configurations of ROCKET-2 in table 5. During training, each trajectory is divided into segments of length 128 to reduce memory requirements. We initialize the view backbone that is used to encode RGB images with DINO weights and freeze it for training efficiency. During inference, ROCKET-2 can access up to 128 key-value attention caches of past observations. Most training parameters follow those from prior works such as Cai et al. (2025a); Baker et al. (2022).
@@
Appendix C Training Datasets
We train ROCKET-2 using 1.6 billion frames of human gameplay data collected by OpenAI (Baker et al., 2022). Each frame is annotated with fine-grained interaction events, including actions such as mining, crafting, item usage, and combat. This large-scale dataset provides rich behavioral diversity across a wide range of tasks and environmental contexts. Building on the protocol established by ROCKET-1 (Cai et al., 2025a), we further enrich each interaction event with frame-level object masks that identify the involved entities or objects. For each interaction event, we ensure that all associated masks correspond to the same target entity. From these, one masked frame is randomly sampled to serve as the cross-view goal specification for training. This process enables the model to learn to align its current egocentric observation with high-level goals presented from different viewpoints. The interaction types used in training include use, break, approach, craft, and kill entity, following the taxonomy introduced in ROCKET-1. These categories cover a broad spectrum of interaction patterns, equipping ROCKET-2 with the ability to generalize across diverse task structures and object affordances.
Figure 11: Minecraft Interaction Benchmark. The benchmark in ROCKET-1 is designed to test an agent’s interaction capabilities with visual goals. The 12 tasks challenge the agent’s ability to follow visual cues to interact with objects at precise spatial locations, a key advancement over previous benchmarks. Image credit: Cai et al. (2025a).
Appendix D Minecraft Interaction Benchmark
To assess ROCKET-2’s ability to accurately interact with target objects under cross-view settings, we adopt the Minecraft Interaction Benchmark introduced in ROCKET-1. This benchmark consists of 12 tasks grouped into six categories: Hunt, Mine, Interact, Navigate, Tool, and Place. Each task requires the agent to identify and interact with specific objects or locations based on visual-temporal prompts. An overview of the benchmark is shown in fig. 11. Compared to prior benchmarks such as Lin et al. (2023), this benchmark places a stronger emphasis on precise spatial interaction—requiring the agent not just to recognize object categories, but to distinguish between multiple instances of the same object type and interact with the correct one at a particular location. This setup significantly raises the bar for visual grounding and spatial understanding, making it a more rigorous testbed for evaluating cross-view interaction capabilities in partially observable 3D environments.
Appendix E Analyzing Failure Cases
We conduct a detailed analysis of failure cases during ROCKET-2’s task execution and identify three primary sources of error:
Prediction Drift
When pursuing distant targets over extended sequences, the predicted interaction point gradually drifts away from the actual object. This issue arises because the agent relies heavily on temporal consistency from memory to maintain object identity. However, during training, the model was only exposed to sequences of up to 128 steps. As a result, it struggles to generalize to longer horizons at test time, revealing limitations in its long-range memory modeling.
Distance Perception Error
In cases where the agent’s current egocentric view differs significantly from the goal-provided view, we observe that the agent sometimes stops one step short of the target, leading to failed interactions. This suggests a misjudgment of spatial proximity under high viewpoint discrepancy. In particular, updating the goal to reflect the current view of the agent mitigates this problem, indicating that the model encounters larger cross-view gaps during inference than during training, affecting its precision of spatial reasoning.
Action Jitter
When evaluating the original version of ROCKET-2, we notice substantial action jitter—rapid, unstable changes in predicted actions—which often results in failures for tasks requiring fine control, such as block placement. We find that feeding previous actions as input during both training and inference significantly improves action stability, leading to smoother and more reliable behavior.
Appendix F Comparison with ROCKET-1
ROCKET-1 tackles interaction problems in 3D worlds by training a visuomotor policy to identify interaction targets based on semantic segmentations in the visual context. While it resolves traditional goal images’ ambiguity and generation challenges, it relies on SAM-2 (Ravi et al., 2024) to track goals and segment them during inference, severely limiting real-time performance. Additionally, its frame-by-frame segmentation training fails to build the policy’s 3D spatial perception, making it heavily dependent on SAM-2’s tracking capabilities. ROCKET-2 is a cross-view segmentation-conditioned policy that. Unlike ROCKET-1, it enables the policy to align the goal across camera views by itself and removes the need for real-time segmentation.
Appendix G Hindsight Trajectory Relabeling
There are two main approaches to collecting labeled trajectory data: (1) providing instructions for contractors to collect trajectories in real-time, ensuring a causal link between actions and labels but incurring high costs and limited scalability; and (2) gathering large amounts of trajectories and generating labels through post-processing, known as hindsight trajectory relabeling. While the first approach (Lynch et al., 2023; Padalkar et al., 2023) produces higher-quality data, its cost constraints have led most research to adopt the second, more scalable one. Andrychowicz et al. (2017) was the first to reinterpret a trajectory’s behavior using its final frame, greatly improving data utilization and inspiring subsequent research on goal-conditioned policies. Lifshitz et al. (2023) extended the approach by using the last 16 frames of a trajectory as the more expressive goal. Sundaresan et al. introduced hand-drawn sketches as a goal modality, reducing semantic ambiguity. Gu et al. (2023) converted the robotic end-effector moving sketch to a 2D image as the goal, providing richer procedural details. ROCKET-1 introduced backward trajectory relabeling, which first identifies interaction objects and events, then utilizes object tracking models (Ravi et al., 2024) to generate frame-level segmentations. This data supports training segmentation-conditioned policies. Our paper explores using this dataset to train policies for cross-view goal alignment.
@@ Appendix G Hindsight Trajectory Relabeling
Table 6: ROCKETs Inference Pipeline Formulation. Molmo can locate the target object based on the task prompt. SAM uses the point to generate object mask $m_{t}$ w.r.t. $o_{t}$ and supports per-frame tracking.
@@
Figure 12: Zero-Shot Demonstrations. We evaluate ROCKET-2 in a variety of previously unseen environments, including Unreal Engine (Zhong et al., 2024), Doom (Kempka et al., 2016), DeepMind Lab (Beattie et al., 2016), Portal, and Dark Souls. Remarkably, ROCKET-2 exhibits strong exploration abilities guided by visual cues in Unreal Engine, successfully transfers certain skills such as attacking in Doom, and maintains robustness to observations that differ significantly from those in Minecraft (Unreal Engine, Portal, and Dark Souls).
Figure 13: Training Dataset. Each row represents a distinct interaction event (such as use, break, or approach) in which all segmentation masks correspond to the same goal. The left column presents sampled cross-view goals and their associated masks for each event, while the right column displays the corresponding frames during that event. This pairing of a cross-view objective with an egocentric action frame constitutes the foundation of our training dataset. In the images, the white point marks the center of the mask, whereas the green point indicates the location predicted by ROCKET-2.